OpenStudy (anonymous):

In the below system, solve for y in the first equation. x – 3y = –6 2x – 7y = 10

OpenStudy (chriss):

x-3y=-6 x+6=3y \[\frac{x+6}{3}= y\] \[\frac{x}{3}+2=y\]